# Suite-IN++: A FlexiWear BodyNet Integrating Global and Local Motion Features from Apple Suite for Robust Inertial Navigation

### v1

@@ Abstract
Index Terms:
Wearable devices, inertial navigation, contrastive learning, golbal and local motion features.I Introduction
@@
With the rapid advancement of mobile computing and wearable technology, smartphones and other wearable devices have become integral to daily life [1, 2, 3, 4, 5]. These devices are increasingly equipped with inertial measurement units (IMUs) that enable comprehensive motion capture through various body attachments [6, 7]. This technological evolution has driven significant progress in human-centric computing applications, including human pose estimation [6, 8, 9, 10], activity recognition [11, 7], and pedestrian localization [5, 12, 13, 12, 14, 15].
@@
Pedestrian dead reckoning (PDR) has attracted considerable attention for its ability to provide continuous and flexible positioning [16, 12, 17, 18]. It utilizes IMU data along with biomechanical constraints, such as zero-velocity updates [19, 20] and step-length priors [21, 22, 23], to estimate movement trajectories in GNSS-denied environments. Recent advances in deep learning have enabled data-driven approaches [12, 24] that extract high-level motion representations, significantly improving positioning accuracy[24, 4, 25, 12, 26]. However, relying on a single sensor, such as a smartphone, for pedestrian localization makes it challenging to robustly accommodate the diverse motion modes of pedestrians in real-life scenarios.
@@
As wearable technology becomes more widespread, utilizing multiple wearable devices to form a sensor network has emerged as a promising approach for human motion estimation [27, 2], with the potential to enhance localization performance across diverse motion modes [28, 29, 30, 31]. Modern users typically carry multiple smart devices (e.g., smartphones, smartwatches, and headphones) that can collectively form a body-area sensor network [32, 9, 33, 34]. However, in real-life scenarios, natural changes in device attachment and the complexity of walking modes make the body network structure more intricate, which we define as the flexiwear bodynet (flexible wearable body network). In a flexiwear bodynet, devices worn on different body parts (e.g., wrist, ears) capture complementary motion characteristics [7, 6]: global motion features describe torso-level displacement and orientation, while local features encode limb-specific dynamics. For pedestrian localization, global features provide essential step direction and velocity estimates, whereas local features offer motion-specific details that can mitigate errors from irregular motions [35]. Effectively integrating these complementary information sources remains a fundamental challenge in multi-device positioning systems, especially in complex real-life scenarios.
@@
Our key insight is that wearable sensors positioned at different body parts inherently capture both global and local motion characteristics. In pedestrian localization, these two feature types play distinct roles: Global features (torso displacement, rotation, and velocity) dominate trajectory estimation, while local features (arm swing amplitude, step cadence, etc.) provide motion-specific refinements. Our previous work Suite-IN [3] demonstrated the potential of aggregating global motion features from multiple devices, but reveals two critical limitations: (1) Inadequate Global Feature Aggregation: The arithmetic mean fusion of global features fails to account for device-specific information quality variations caused by body placement differences. (2) Underutilized Local Features: The contrastive learning framework treats local motion as mere noise, disregarding its potential for enhancing position estimation.
@@
To overcome these limitations, we extend Suite-IN [3] and introduce Suite-IN++ shown in Fig. 1, which introduces two key improvements: First, instead of averaging global features, we adopt a weighted fusion strategy that evaluates the reliability of each device’s motion information, enabling a more informed and adaptive aggregation of global motion features. Second, rather than disregarding local motion as noise, we incorporate an attention mechanism to capture intrinsic relationships between different local motion features. This allows our model to extract motion details—such as gait changes and walking modes—that enhance positioning accuracy and generalization. In addition, unlike previous work [36, 37, 38], we use real-life wearable devices (Apple Suite: iPhone, Apple Watch and Airpods) for data collection. The IMUs in these consumer-grade devices tend to have higher noise levels compared to specialized IMU systems like Xsens [39] and Noitom [40], introducing additional challenges for accurate pedestrian localization.
@@ II-A Data-driven Pedestrian Localization
Data-driven method are proposed to directly estimate position from IMU data. Data-driven smartphone inertial odometry has gained interest in recent years. IONet [24] is a neural network-based inertial navigation method, which uses a long short-term memory (LSTM) network model to regress pedestrian velocity magnitude and the rate of motion-heading change from smartphone data. RIDI[26] classifies the phone attachment by a support vector machine, and then regresses velocity for each attachment. Inertial measurements are distributed differently across domains, motiontransformer [5] exploits generative adversarial network (GAN) and domain adaptation to improve the effectiveness of inertial navigation systems for unseen domains without any paired data. RoNIN[12] uses three different neural network models (LSTM, ResNet, TCN) to achieve end-to-end pedestrian positioning. Based on RIDI, it extends the application scenarios of the inertial navigation system and supports more walking modes. In order to deploy the model on real-life devices, methods such as IMUNet[41] and L-IONet[42] are committed to developing lightweight networks to optimize network design, reduce network parameters and improve operating efficiency without affecting positioning accuracy.
@@ II-B Multi-sensor Fusion for Wearable Sensors
Multi-sensor fusion is the technique that involves gathering and combining information from multiple sensors in order to provide better information for target regression or recognition. The fusion methods can be divided into either signal, feature or decision level fusion[43].
By combining multi-sensor information through multi-sensor fusion, the performance of various wearable sensors-based tasks can be enhanced[37, 36, 32, 44, 45]. Gong et al.[37]proposes a multi-sensor fusion pipeline called DeepIT, which integrates IMU measurements of smartphones and associated earbuds through a reliability network to achieve inertial tracking. Restrained-Weighted-Fusion is introduced by Song et al.[36] to enhance fusion accuracy and robustness of multi-node fusion positioning, and gumbel softmax resampling is used to optimize the weight of each sensor. Sensehar[32] improves human activity recognition performance by extracting features shared by multiple sensors to represent multi-sensor fusion feature. DeepSense[45] is a unified model to fuse multiple similar sensors to solve both classification and regression problems such as human activity recognition and motion tracking. DeepFusion[44] further considers the fusion weights and cross-sensor correlations of different sensors, complementarily utilizing multi-sensor information.
@@ II-C Contrastive Learning for Wearable Sensors
Contrastive learning [46, 47, 48, 11, 49] has been widely explored and applied in wearable-based human-centric tasks. The core principle of contrastive learning is to acquire robust representations by distinguishing between similar and dissimilar instances, often optimized using InfoNCE loss or its variants [46]. For instance, COCOA [46] leverages contrastive learning to extract high-quality representations from multi-sensor data by computing cross-correlations between different modalities while minimizing the similarity between unrelated instances. Recognizing the importance of modality-specific features in downstream tasks, Liu et al. [47] introduce an orthogonality constraint, enabling the simultaneous utilization of both modality-shared and modality-specific representations through contrastive loss. For target modality data that lacks label, learning from the best[48] ultilizes the contrastive representation misalignment loss between the source and target modality to extract the share feature of two modalities. The contrastive learning in weakly supervised settings provides the supervision for unlabeled data, which bridges the gap between human activity classification and segmentation tasks. Xia et al. [11] adopts the sample-to-prototype contrast module for further refining the rough activity recognition results (recognition task) in the sequence to the prediction of each sample’s activity (segmentation task).
@@ III-A Dataset Overview
Our dataset comprises 429 sequences, totaling approximately 20 hours of recordings and 54.5 km of walking distance, collected from 12 participants across 14 different scenes spanning two cities and four buildings. To simulate realistic usage conditions, participants were instructed to wear devices in various configurations (e.g., handheld, in-pocket, in-backpack) and perform diverse daily walking behaviors (e.g., removing a device, sitting down, standing still, squatting mid-walk). Compared with existing mainstream data sets such as OxIOD[50], RoNIN[12] and DeepIT[37], our dataset is significantly larger, covers a broader range of walking areas and device configurations, and captures more complex motion patterns. The comprehensive dataset statistics are presented in Tab. I, where analysis of the RoNIN dataset is restricted to its publicly accessible portion due to limited data availability constraints
To capture varying levels of motion complexity, we define five walking modes, summarized in Tab. II. The simplest mode, STW, involves stable and consistent walking, while subsequent modes (PVW, MVW) introduce motion disturbances to one or multiple devices. The most challenging modes (DRW, DLW) simulate device removal and incorporate everyday activities. This progressive structure makes the dataset well-suited for modeling real-life pedestrian motion. The diversity in device configurations and walking modes also makes it applicable beyond localization, including tasks such as human action recognition [11, 45, 44] and pose estimation [6, 9, 10].
@@ III-B Data Collection Apparatus
To obtain ground-truth trajectories, we used ARKit 333https://developer.apple.com/documentation/arkit/, a tightly-coupled filtering-based visual-inertial odometry (VIO) framework integrated into iOS. ARKit achieves a drift error of approximately 0.02 m per second [51]. We use a harness to attach a 3D tracking phone (iPhone 11) to a body to obtain the ground truth position frome ARKit and let subjects handle the other phone freely for IMU data collection. The 3D tracking phone records pose estimates at 30 Hz, represented by a translation vector and a unit quaternion. Ground-truth is only collected for the 3D tracking phone attached to a harness, as our objective is to estimate the human body’s trajectory rather than that of the device movement. Fig. 2 shows how the devices are worn and the specific settings of the dataset: flexible device configuration and complex walking modes.
@@ III-C Data Processing
Before the data-collection, we performed sensor bias calibration and spatial alignment between the data-collection phone and tracking phone. The rotation matrix at the initial timestamp was used to project IMU readings from the phone into the global coordinate system defined by VIO. The smartwatch and headphones were natively aligned with the data-collection phone during acquisition. Accurate temporal synchronization between VIO and IMU is crucial due to the high-frequency and time-sensitive nature of inertial data. Following the protocol from [52], participants were instructed to jump three times at the beginning and end of each session. These sharp vertical spikes were used to align timestamps by matching peak patterns between VIO and IMU sequences.
@@ IV-B Decouple of global and local motion features for trajectory regression
The independent feature extractors and have the same structure composed of six 1D CNN blocks but do not share weights. In our model, the activation function ReLU, batch normalization and dropout technique are leveraged in CNN blocks, with max pooling along the temporal and sensor channel dimension to extract motion features from the wearable devices. It is worth noting that the convolution kernel size is along the temporal dimension, ensuring that the convolution slides over the time dimension to capture temporal patterns without altering the sensor channel dimensionality [7, 53].
Contrastive learning for global and local motion feature separation: In complex walking modes, global features capture the overall movement trend, while local features reflect subtle dynamics. Distinguishing between them is crucial for creating highly discriminative motion representations. Contrastive learning uses the positive and negative sample pairs to enable the model to automatically strengthen the difference between the two types of features in self-supervised training, thereby effectively distinguishing and capturing global and local information. For the same sensory data, our model iterates over all extracted features, regarding global features of different modalities as positive pairs, regarding global feature and local feature of each modality as negative pairs, and regarding local features between different modalities as negative pairs. We calculate the constractive loss following InfoNCE loss[54, 47] with these positive and negative pairs, represesnted by:
@@
Orthogonality constraint for global and local space: To prevent global motion information from contaminating the local feature space and to ensure that each wearable device captures its unique local motion patterns, we introduce orthogonality constraints. Drawing inspiration from [47, 55], we enforce these constraints both between the global and local features within the same modality and among the local features across different modalities. This design ensures that the decomposed global and local feature space captures independent semantic information, with each feature subset contributing uniquely to motion representation. To implement these constraints, we minimize the angular similarities between the corresponding feature pairs using a cosine embedding loss, which can be expressed as:
Velocity and trajectory regression: After extracting global and local motion features from each wearable device, the global motion features can be aggregated to estimate the direction and velocity of the human body. A common aggregation method is to take the arithmetic average () of the global motion features of all devices[32, 3]. The aggregated shared features are represented as , which fuses multi-device data into a shared low-dimensional latent space to represent the aggregated global motion features.
We take the average velocity of the window as the network’s output inspired by previous works[36, 12]. A fully connected layer is used to simply process the aggregated global motion features to obtain the global movement velocity :
The global movement velocity is supervised by the Mean-Squared-Error (MSE)[56] loss:
@@
The flexiwear bodynet presents two key challenges for pedestrian localization: (1) flexible device configurations and (2) complex walking modes. Relying solely on global motion features proves insufficient under these conditions, as shown in the limitations of our previous work, Suite-IN[3]: (1) Inadequate Global Feature Aggregation: Suite-IN employs simple arithmetic averaging to aggregate global features, disregarding the varying reliability of motion information across devices. Due to the differences in hardware quality and placement on the body, the amount of reliable global motion information carried by different devices varies. For instance, during walking, motion information captured by headphones worn on the head tends to be more stable in reflecting overall displacement. Under flexible device configurations, device attachment may change dynamically, causing the reliability of motion information from each device to fluctuate. Therefore, aggregating global features according to the reliability of each device’s motion information is crucial for capturing overall motion trends and enhancing localization robustness. (2) Underutilized Local Features: Suite-IN treats local features as noise, overlooking their positive contribution to motion estimation. Local features encode fine-grained motion details at each device location, such as wrist swing amplitude and rhythm, which can refine motion estimates, particularly under complex walking modes. Effectively utilizing these local features holds the potential to improve both localization accuracy and system stability.

### v2

@@ Abstract
I Introduction
@@
With the rapid advancement of mobile computing and wearable technology, smartphones and other wearable devices have become integral to daily life [9628113, 9673682, sun2024suite, chen2019deep, chen2019motiontransformer]. These devices are increasingly equipped with inertial measurement units (IMUs) that enable comprehensive motion capture through various body attachments [zhang2024dynamic, pei2021mars]. This technological evolution has driven significant progress in human-centric computing applications, including human pose estimation [zhang2024dynamic, yi2021transpose, mollyn2023imuposer, 9737726], activity recognition [xia2024timestamp, pei2021mars], and pedestrian localization [chen2019motiontransformer, herath2020ronin, wang2022inertial, herath2020ronin, 10035881, jiang2024robust].
@@
Pedestrian dead reckoning (PDR) has attracted considerable attention for its ability to provide continuous and flexible positioning [9406816, herath2020ronin, pei2018optimal, 10478937]. It utilizes IMU data along with biomechanical constraints, such as zero-velocity updates [10535289, 10648849] and step-length priors [10371336, 9511443, 10829797], to estimate movement trajectories in GNSS-denied environments. Recent advances in deep learning have enabled data-driven approaches [herath2020ronin, chen2018ionet] that extract high-level motion representations, significantly improving positioning accuracy[chen2018ionet, chen2019deep, 8937008, herath2020ronin, yan2018ridi]. However, relying on a single sensor, such as a smartphone, for pedestrian localization makes it challenging to robustly accommodate the diverse motion modes of pedestrians in real-life scenarios.
@@
As wearable technology becomes more widespread, utilizing multiple wearable devices to form a sensor network has emerged as a promising approach for human motion estimation [9839489, 9673682], with the potential to enhance localization performance across diverse motion modes [9389629, 9442861, 9841001, 10288089]. Modern users typically carry multiple smart devices (e.g., smartphones, smartwatches, and headphones) that can collectively form a body-area sensor network [jeyakumar2019sensehar, mollyn2023imuposer, 10592803, 10319091]. However, in real-life scenarios, natural changes in device attachment and the complexity of walking modes make the body network structure more intricate, which we define as the flexiwear bodynet (flexible wearable body network). In a flexiwear bodynet, devices worn on different body parts (e.g., wrist, ears) capture complementary motion characteristics [pei2021mars, zhang2024dynamic]: global motion features describe torso-level displacement and orientation, while local features encode limb-specific dynamics. For pedestrian localization, global features provide essential step direction and velocity estimates, whereas local features offer motion-specific details that can mitigate errors from irregular motions [8868212]. Effectively integrating these complementary information sources remains a fundamental challenge in multi-device positioning systems, especially in complex real-life scenarios.
@@
Our key insight is that wearable sensors positioned at different body parts inherently capture both global and local motion characteristics. In pedestrian localization, these two feature types play distinct roles: Global features (torso displacement, rotation, and velocity) dominate trajectory estimation, while local features (arm swing amplitude, step cadence, etc.) provide motion-specific refinements. Our previous work Suite-IN [sun2024suite] demonstrated the potential of aggregating global motion features from multiple devices, but reveals two critical limitations: (1) Inadequate Global Feature Aggregation: The arithmetic mean fusion of global features fails to account for device-specific information quality variations caused by body placement differences. (2) Underutilized Local Features: The contrastive learning framework treats local motion as mere noise, disregarding its potential for enhancing position estimation.
@@
To overcome these limitations, we extend Suite-IN [sun2024suite] and introduce Suite-IN++ shown in Fig. 1, which introduces two key improvements: First, instead of averaging global features, we adopt a weighted fusion strategy that evaluates the reliability of each device’s motion information, enabling a more informed and adaptive aggregation of global motion features. Second, rather than disregarding local motion as noise, we incorporate an attention mechanism to capture intrinsic relationships between different local motion features. This allows our model to extract motion details—such as gait changes and walking modes—that enhance positioning accuracy and generalization. In addition, unlike previous work [10447298, gong2021robust, 8865667], we use real-life wearable devices (Apple Suite: iPhone, Apple Watch and Airpods) for data collection. The IMUs in these consumer-grade devices tend to have higher noise levels compared to specialized IMU systems like Xsens [schepers2018xsens] and Noitom [noitom2019], introducing additional challenges for accurate pedestrian localization.
@@ II-A Data-driven Pedestrian Localization
Data-driven method are proposed to directly estimate position from IMU data. Data-driven smartphone inertial odometry has gained interest in recent years. IONet [chen2018ionet] is a neural network-based inertial navigation method, which uses a long short-term memory (LSTM) network model to regress pedestrian velocity magnitude and the rate of motion-heading change from smartphone data. RIDI[yan2018ridi] classifies the phone attachment by a support vector machine, and then regresses velocity for each attachment. Inertial measurements are distributed differently across domains, motiontransformer [chen2019motiontransformer] exploits generative adversarial network (GAN) and domain adaptation to improve the effectiveness of inertial navigation systems for unseen domains without any paired data. RoNIN[herath2020ronin] uses three different neural network models (LSTM, ResNet, TCN) to achieve end-to-end pedestrian positioning. Based on RIDI, it extends the application scenarios of the inertial navigation system and supports more walking modes. In order to deploy the model on real-life devices, methods such as IMUNet[zeinali2024imunet] and L-IONet[8960327] are committed to developing lightweight networks to optimize network design, reduce network parameters and improve operating efficiency without affecting positioning accuracy.
Beyond existing data-driven localization methods, recent works have focused on signal modeling and enhancement to improve feature robustness and generalization. TG-FE[wang2025general] encodes inertial signals into a memory-based graph and fuses local and global information to extract transferable features. HEROS-GAN[wang2025heros] enhances low-cost accelerometer signals into high-quality equivalents using optimal transport supervision and energy modulation. These approaches offer promising directions for improving inertial localization in real-world scenarios.
@@ II-B Multi-sensor Fusion for Wearable Sensors
As wearable technology advances, there is an opportunity to leverage a variety of smart devices to enhance position estimation and broaden the applicable scenarios[gong2021robust, 10447298]. Multi-sensor fusion is the technique that involves gathering and combining information from multiple sensors in order to provide better information for target regression or recognition. The fusion methods can be divided into either signal, feature or decision level fusion[gravina2017multi].
By combining multi-sensor information through multi-sensor fusion, the performance of various wearable sensors-based tasks can be enhanced[gong2021robust, 10447298, jeyakumar2019sensehar, xue2019deepfusion, yao2017deepsense]. Gong et al.[gong2021robust]proposes a multi-sensor fusion pipeline called DeepIT, which integrates IMU measurements of smartphones and associated earbuds through a reliability network to achieve inertial tracking. Restrained-Weighted-Fusion is introduced by Song et al.[10447298] to enhance fusion accuracy and robustness of multi-node fusion positioning, and gumbel softmax resampling is used to optimize the weight of each sensor. Sensehar[jeyakumar2019sensehar] improves human activity recognition performance by extracting features shared by multiple sensors to represent multi-sensor fusion feature. DeepSense[yao2017deepsense] is a unified model to fuse multiple similar sensors to solve both classification and regression problems such as human activity recognition and motion tracking. DeepFusion[xue2019deepfusion] further considers the fusion weights and cross-sensor correlations of different sensors, complementarily utilizing multi-sensor information.
@@ II-C Contrastive Learning for Wearable Sensors
Contrastive learning [deldari2022cocoa, liu2024focal, fortes2022learning, xia2024timestamp, 10705896] has been widely explored and applied in wearable-based human-centric tasks. The core principle of contrastive learning is to acquire robust representations by distinguishing between similar and dissimilar instances, often optimized using InfoNCE loss or its variants [deldari2022cocoa]. For instance, COCOA [deldari2022cocoa] leverages contrastive learning to extract high-quality representations from multi-sensor data by computing cross-correlations between different modalities while minimizing the similarity between unrelated instances. Recognizing the importance of modality-specific features in downstream tasks, Liu et al. [liu2024focal] introduce an orthogonality constraint, enabling the simultaneous utilization of both modality-shared and modality-specific representations through contrastive loss. For target modality data that lacks label, learning from the best[fortes2022learning] ultilizes the contrastive representation misalignment loss between the source and target modality to extract the share feature of two modalities. The contrastive learning in weakly supervised settings provides the supervision for unlabeled data, which bridges the gap between human activity classification and segmentation tasks. Xia et al. [xia2024timestamp] adopts the sample-to-prototype contrast module for further refining the rough activity recognition results (recognition task) in the sequence to the prediction of each sample’s activity (segmentation task).
@@ III-A Dataset Overview
Our dataset comprises 429 sequences, totaling approximately 20 hours of recordings and 54.5 km of walking distance, collected from 12 participants across 14 different scenes spanning two cities and four buildings. To simulate realistic usage conditions, participants were instructed to wear devices in various configurations (e.g., handheld, in-pocket, in-backpack) and perform diverse daily walking behaviors (e.g., removing a device, sitting down, standing still, squatting mid-walk). Compared with existing mainstream data sets such as OxIOD[chen2020deep], RoNIN[herath2020ronin] and DeepIT[gong2021robust], our dataset is significantly larger, covers a broader range of walking areas and device configurations, and captures more complex motion patterns. The comprehensive dataset statistics are presented in Tab. I, where analysis of the RoNIN dataset is restricted to its publicly accessible portion due to limited data availability constraints
To capture varying levels of motion complexity, we define five walking modes, summarized in Tab. II. The simplest mode, STW, involves stable and consistent walking, while subsequent modes (PVW, MVW) introduce motion disturbances to one or multiple devices. The most challenging modes (DRW, DLW) simulate device removal and incorporate everyday activities. This progressive structure makes the dataset well-suited for modeling real-life pedestrian motion. The diversity in device configurations and walking modes also makes it applicable beyond localization, including tasks such as human action recognition [xia2024timestamp, yao2017deepsense, xue2019deepfusion] and pose estimation [zhang2024dynamic, mollyn2023imuposer, 9737726].
@@ III-B Data Collection Apparatus
To obtain ground-truth trajectories, we used ARKit 333https://developer.apple.com/documentation/arkit/, a tightly-coupled filtering-based visual-inertial odometry (VIO) framework integrated into iOS. ARKit achieves a drift error of approximately 0.02 m per second [kim2022benchmark]. We use a harness to attach a 3D tracking phone (iPhone 11) to a body to obtain the ground truth position frome ARKit and let subjects handle the other phone freely for IMU data collection. The 3D tracking phone records pose estimates at 30 Hz, represented by a translation vector and a unit quaternion. Ground-truth is only collected for the 3D tracking phone attached to a harness, as our objective is to estimate the human body’s trajectory rather than that of the device movement. Fig. 2 shows how the devices are worn and the specific settings of the dataset: flexible device configuration and complex walking modes.
@@ III-C Data Processing
Before the data-collection, we performed sensor bias calibration and spatial alignment between the data-collection phone and tracking phone. The rotation matrix at the initial timestamp was used to project IMU readings from the phone into the global coordinate system defined by VIO. The smartwatch and headphones were natively aligned with the data-collection phone during acquisition. Accurate temporal synchronization between VIO and IMU is crucial due to the high-frequency and time-sensitive nature of inertial data. Following the protocol from [li2021motion], participants were instructed to jump three times at the beginning and end of each session. These sharp vertical spikes were used to align timestamps by matching peak patterns between VIO and IMU sequences.
@@ IV-B Decouple of global and local motion features for trajectory regression
The independent feature extractors and have the same structure composed of six 1D CNN blocks but do not share weights. In our model, the activation function ReLU, batch normalization and dropout technique are leveraged in CNN blocks, with max pooling along the temporal and sensor channel dimension to extract motion features from the wearable devices. It is worth noting that the convolution kernel size is along the temporal dimension, ensuring that the convolution slides over the time dimension to capture temporal patterns without altering the sensor channel dimensionality [pei2021mars, 10899893].
Contrastive learning for global and local motion feature separation: In complex walking modes, global features capture the overall movement trend, while local features reflect subtle dynamics. Distinguishing between them is crucial for creating highly discriminative motion representations. Contrastive learning uses the positive and negative sample pairs to enable the model to automatically strengthen the difference between the two types of features in self-supervised training, thereby effectively distinguishing and capturing global and local information. For the same sensory data, our model iterates over all extracted features, regarding global features of different modalities as positive pairs, regarding global feature and local feature of each modality as negative pairs, and regarding local features between different modalities as negative pairs. We calculate the constractive loss following InfoNCE loss[tian2020contrastive, liu2024focal] with these positive and negative pairs, represesnted by:
@@
Orthogonality constraint for global and local space: To prevent global motion information from contaminating the local feature space and to ensure that each wearable device captures its unique local motion patterns, we introduce orthogonality constraints. Drawing inspiration from [liu2024focal, cai2024orthogonality], we enforce these constraints both between the global and local features within the same modality and among the local features across different modalities. This design ensures that the decomposed global and local feature space captures independent semantic information, with each feature subset contributing uniquely to motion representation. To implement these constraints, we minimize the angular similarities between the corresponding feature pairs using a cosine embedding loss, which can be expressed as:
Velocity and trajectory regression: After extracting global and local motion features from each wearable device, the global motion features can be aggregated to estimate the direction and velocity of the human body. A common aggregation method is to take the arithmetic average () of the global motion features of all devices[jeyakumar2019sensehar, sun2024suite]. The aggregated shared features are represented as , which fuses multi-device data into a shared low-dimensional latent space to represent the aggregated global motion features.
We take the average velocity of the window as the network’s output inspired by previous works[10447298, herath2020ronin]. A fully connected layer is used to simply process the aggregated global motion features to obtain the global movement velocity :
The global movement velocity is supervised by the Mean-Squared-Error (MSE)[koksoy2006multiresponse] loss:
@@
The flexiwear bodynet presents two key challenges for pedestrian localization: (1) flexible device configurations and (2) complex walking modes. Relying solely on global motion features proves insufficient under these conditions, as shown in the limitations of our previous work, Suite-IN[sun2024suite]: (1) Inadequate Global Feature Aggregation: Suite-IN employs simple arithmetic averaging to aggregate global features, disregarding the varying reliability of motion information across devices. Due to the differences in hardware quality and placement on the body, the amount of reliable global motion information carried by different devices varies. For instance, during walking, motion information captured by headphones worn on the head tends to be more stable in reflecting overall displacement. Under flexible device configurations, device attachment may change dynamically, causing the reliability of motion information from each device to fluctuate. Therefore, aggregating global features according to the reliability of each device’s motion information is crucial for capturing overall motion trends and enhancing localization robustness. (2) Underutilized Local Features: Suite-IN treats local features as noise, overlooking their positive contribution to motion estimation. Local features encode fine-grained motion details at each device location, such as wrist swing amplitude and rhythm, which can refine motion estimates, particularly under complex walking modes. Effectively utilizing these local features holds the potential to improve both localization accuracy and system stability.